My son brought this problem home to me. I'm not totally positive this is the way to solve. Can you confirm?
63-|-78|=?
I told him he had to change the -78 to a positive then follow through the problem.
63-78= -15
Is this correct?

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

63-|-78| = 63-(78)

63-|-78| = 63-78

63-|-78| = -15

So you are correct.

Answer by stanbon(75887)

63-|-78|=?
I told him he had to change the -78 to a positive then follow through the problem.
63-78= -15
------
You are correct.
Cheers,
Stan H.